# Better Aligned with Survey Respondents or Training Data? Unveiling Political Leanings of LLMs on U.S. Supreme Court Cases

### v2

@@ 4.2 Measuring the LLMs Alignments
4.3 Testing for Significance of Stronger Alignments
@@
Figure 4: Pearson Alignment. Cell $(i,j)$ represents the Pearson correlation $\rho$ of LLM $i$ to entity $j$. * shows p-value < 0.05, **∗ ∗ shows p-value < 0.001.
@@ 5 Results and Anyalsis
Furthermore, all LLMs, with the exception of Bloomz and T0, display a significant positive correlation with the three training corpora: Dolma, RedPajama, and Oscar. This alignment may stem from the similar political leanings in these corpora and the models trained on them [^9] Fig 11 in App F presents the alignments between different training corpora and surveyed human opinions. . In contrast, our findings indicate that there are generally no significant alignments between the LLMs’ outputs and surveyed human opinions. The only LLMs that do not follow this trend are LLama3-70b and T0, which we will discuss further in Takeaway 3.
@@
Similarly, the Pearson correlation results in Fig 4 indicate that T0 exhibits a significant correlation only with $E_{\text{pub}}$ (surveyed human opinions), while no significant correlations are observed with other entities. At first glance, this might suggest that T0 is most aligned with human opinions among all entities. However, the significance test results in Fig 5(e) reveal inconsistencies. While correlation between $(T0,E_{\text{pub}})$ is significantly stronger than the correlation between $(T0,E_{\text{court}})$, no such significant differences are found with other entities, such as with $E_{\text{dem}}$ or any of the training corpora. This means that we can only conclude that T0 aligns more closely with surveyed human opinions $E_{\text{pub}}$ than with the court $E_{\text{court}}$ , but we cannot determine whether its alignment with $E_{\text{pub}}$ is significantly stronger than its alignment with other entities, even though there are great differences in $\rho$-values observed in Fig 4.
@@ 6 Implications and Future Directions
Our research reveals that most LLMs exhibit alignment with their training corpora, yet not necessarily with the surveyed human opinions. Nonetheless, in the public discourse framework, attributing human characteristics to AI - known as anthropomorphizing — seems to be quite natural. This tendency may lead to an over-reliance on AI, as users might confuse AI-generated responses for human beings, leading to excessive trust PAIR (2019). Further, anthropomorphism can obscure accountability, shifting the responsibility away from developers and onto the LLM itself. Recent studies McCoy et al. (2024) suggest moving away from anthropomorphic and advocate for a reframing of public dialogue in alternative conceptual frameworks, such as viewing LLMs as simulation systems of the integration of diverse perspectives in their training data Shanahan et al. (2023). In conclusion, fostering a clear public understanding of the distinctions between AI and human beings is essential for a more responsible engagement with AI technologies.
@@ 7 Conclusion
We introduced a pipeline to investigate political leaning in the pretraining corpora, which allows us to compare the LLMs’ political leaning not only with surveyed human opinions but also the political leanings embedded in their pretraining corpora. By examining LLMs’ stances on political issues derived from U.S. Supreme Court cases, our results reveal a significant alignment between the models and their training corpora, yet no similarly strong alignment with human opinions as survyed is found. We call on the AI community to explore methods for rigorous evaluation of LLM alignment, and advocate for more transparent and collaborative strategies in curating training data for LLM.
@@ Acknowledgments
We thank Rashid Haddad, Lukas Gosch, Valentin Hoffman, Max Prior and Yang Li for the helpful discussions.
@@ Appendix E Quality Assessment of Stance Detection
Figure 10: Distribution of length of all the matched documents.
Appendix F Corpora-Human Alignment
Appendix G Post-training
The key difference between OLMo-SFT (Supervised Fine-Tuning) and OLMo-Instruct lies in their fine-tuning objectives and intended uses. OLMo-SFT is fine-tuned for general language tasks using labeled data, using the TULU dataset Ivison et al. (2023). It is optimized for structured responses but isn’t specifically trained to follow user instructions. OLMo-instruct is further fine-tuned to follow human instructions, using the Ultrafeedback dataset Cui et al. (2023). It is optimized for handling detailed user instructions and conversational prompts, ideal for interactive and task-oriented use.
Appendix H Williams Test Results
Appendix I Example of a Retrieved Document

### v3

@@ 4.2 Measuring the LLMs Alignments
4.3 Testing for Significance of Alignments
@@
Figure 5: Pearson Alignment. Cell $(i,j)$ represents the Pearson correlation $\rho$ of LLM $i$ to entity $j$. * shows p-value < 0.05, **∗ ∗ shows p-value < 0.001. random_1 stands for randomized values used as a baseline.
@@ 5 Results and Anyalsis
Furthermore, all LLMs, with the exception of Bloomz and T0, display a significant positive correlation with the three training corpora: Dolma, RedPajama, and Oscar. This alignment may stem from the similar political leanings in these corpora and the models trained on them. [^9] Fig 12 in App F presents the alignments between different training corpora and surveyed human opinions. In contrast, our findings indicate that there are generally no significant alignments between the LLMs’ outputs and surveyed human opinions. The only LLMs that do not follow this trend are LLama3-70b and T0, which we will discuss further in Takeaway 3.
@@
Similarly, the Pearson correlation results in Fig 5 indicate that T0 exhibits a significant correlation only with $E_{\text{pub}}$ (surveyed human opinions), while no significant correlations are observed with other entities. At first glance, this might suggest that T0 is most aligned with human opinions among all entities. However, the significance test results in the Subfig (e) in Fig 6 reveal inconsistencies. While correlation between $(T0,E_{\text{pub}})$ is significantly stronger than the correlation between $(T0,E_{\text{court}})$, no such significant differences are found with other entities, such as with $E_{\text{dem}}$ or any of the training corpora. This means that we can only conclude that T0 aligns more closely with surveyed human opinions $E_{\text{pub}}$ than with the court $E_{\text{court}}$ , but we cannot determine whether its alignment with $E_{\text{pub}}$ is significantly stronger than its alignment with other entities, even though there are great differences in $\rho$-values observed in Fig 5.
@@ 6 Implications and Future Directions
Our research reveals that most LLMs exhibit alignment with their training corpora, yet not necessarily with the surveyed human opinions. Nonetheless, in the public discourse framework, attributing human characteristics to AI, also known as anthropomorphizing, seems to be quite natural. This tendency may lead to an over-reliance on AI, as users might confuse AI-generated responses for human beings, leading to excessive trust PAIR (2019). Further, anthropomorphism can obscure accountability, shifting the responsibility away from developers and onto the LLM itself. Recent studies McCoy et al. (2024) suggest moving away from anthropomorphic and advocate for a reframing of public dialogue in alternative conceptual frameworks, such as viewing LLMs as simulation systems of the integration of diverse perspectives in their training data Shanahan et al. (2023). In conclusion, fostering a clear public understanding of the distinctions between AI and human beings is essential for a more responsible engagement with AI technologies.
@@ 7 Conclusion
We introduced a pipeline to investigate political leaning in the pretraining corpora, which allows us to compare the LLMs’ political leaning not only with surveyed human opinions but also the political leanings embedded in their pretraining corpora. By examining LLMs’ stances on political issues derived from U.S. Supreme Court cases, our results reveal a significant alignment between the models and their training corpora, yet no similarly strong alignment with human opinions is found. These findings suggest that political bias in LLMs may be at least partly a result of memorization of biased content from pretraining corpora. We call on the AI community to explore methods for detecting, and mitigating memorized political bias in LLMs, and advocate for more transparent and collaborative strategies in curating training data for LLMs.
@@ Acknowledgments
We thank Rashid Haddad, Lukas Gosch, Valentin Hoffman, Max Prior and Yang Li for the helpful discussions. BP is supported by the ERC Consolidator Grant DIALECT 101043235.
@@ Appendix E Quality Assessment of Stance Detection
Figure 10: Distribution of length of all the matched documents.
Appendix F Corpora-Human Alignment
Appendix G Post-training
The key difference between OLMo-SFT (Supervised Fine-Tuning) and OLMo-Instruct lies in their fine-tuning objectives and intended uses. OLMo-SFT is fine-tuned for general language tasks using labeled data, using the TULU datasetIvison* et al. (2023). It is optimized for structured responses but isn’t specifically trained to follow user instructions. OLMo-instruct is further fine-tuned to follow human instructions, using the Ultrafeedback dataset Cui et al. (2023). It is optimized for handling detailed user instructions and conversational prompts, ideal for interactive and task-oriented use.
Appendix H Williams Test Results
Appendix I Example of a Retrieved Document